OpenStudy (anonymous):

What are the zeros of the polynomial function f(x) = x3 – 4x2 – 12x

OpenStudy (anonymous):

it's x cubed - 4x squared- 12x

OpenStudy (zpupster):

factor x out x(x^2-4x-12)=0 then factor the quadreatic x(x-6)(x+2)=0 set all to 0 and solve x=0 (x+2)=0 (x-6)=0